Permutation or Combination:
Selecting a president, vice president, and treasurer of the 100-student class.
How do I know if it's permutation or combination?

Answer by jim_thompson5910(35256)

Order is important and order matters. Say you had three people and you called them A, B, C. The order ABC is different from BAC because A is president in the first group and B is president in the second group.


So again, order matters. This means we'll be doing a permutation. You could use the nPr formula, but below is a much simpler way.


You have 100 choices for slot 1 (president), 100-1=99 choices for slot 2 (VP), and 100-2 = 98 choices for slot 3 (treasurer). Notice how as you progress through the slots, you lose one person each time. This is because no person can serve multiple posts.


Multiply these choices out: 100*99*98 = 970,200


So there are 970,200 different ways to do this.